OpenStudy (anonymous):

evaluate the sum of the finite geometric series. 4+12+36+108+...+972

OpenStudy (anonymous):

this is a geometric sequence which r = 3 972 = 4*3^(n-1) n= 6 so sum = 4*(1-3^6)/1-3 = 4*(-728)/-2 = 1456

OpenStudy (anonymous):

\[4(1+3+3^2+3^3+3^4+3^5)=4\times \frac{3^6-1}{3-1}\]

OpenStudy (anonymous):

how does n=6

OpenStudy (anonymous):

do you understand what i wrote?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i understand the equation better, but i need to know how to get N

OpenStudy (anonymous):

which side, the left hand side of the equal sign or the right hand side? the left hand side i got by noting that each term had a common factor of 4, so i factored it out

OpenStudy (anonymous):

that left me with \[4(1+3+9+27+81+243)\] and these are all clearly powers of 3, so i rewrote it as powers of 3 as \[4(1+3+3^2+3^3+3^4+3^5)\]

OpenStudy (anonymous):

my teacher said to use this equation S = a*((1 - r^n)/(1-r))

OpenStudy (anonymous):

then the right hand side i get by knowing that one more than 5 is 6, so it must be \[4\times\frac{3^6-1}{3-1}\] and it is arithmetic from there on in

OpenStudy (anonymous):

you teachers is telling you to use a formula, which is fine, that is what i did but do not fret about what \(n\) or \( n +1\) is. \(n \) is the last power you see, in this case it was 5, so \(n + 1\) is \(5+1\) is 6

OpenStudy (anonymous):

where you get the 5

OpenStudy (anonymous):

look at the left hand side of the equal sign. i got it by computing

OpenStudy (anonymous):

lets go slow do you agree that we can factor out a 4 from each term and get \[4(1+3+9+27+81+243)\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well if you cannot, then the answer is wrong

OpenStudy (anonymous):

why you cant factor a 4 from 9

OpenStudy (anonymous):

you started with this \[4+12+36+108+...+972\] right?

OpenStudy (anonymous):

oh yeah, lol my bad

OpenStudy (anonymous):

so each term you see has a factor of 4 i factored it out of the whole thing. that is the first step

OpenStudy (anonymous):

yeah, i know. I thought you meant 3,9, and 27.

OpenStudy (anonymous):

oh no, i meant starting with what you wrote

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so now again, we have \[4(1+3+9+27+81+243)\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok and each term after the 4 is a power of 3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so lets write it out again, this time as powers of 3 \[4(1+3+9+27+81+243)=4(1+3+3^2+3^3+3^4+3^5)\] is that part ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good. now you tell me, where does the five come from?

OpenStudy (anonymous):

there is 5 terms

OpenStudy (anonymous):

it is right there in the problem yes? it comes from the fact that when you write it in this form, the last term you are looking at is \(3^5\) i see it with my eyes.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hope that answers the question "where does the five come from" it is from computing what you have. then you can use the formula \[4\times \frac{3^6-1}{3-1}=4\times \frac{729-1}{2}=2\times 728=1456\]

OpenStudy (anonymous):

Yes, thanks

OpenStudy (anonymous):

yw